# Reinforcing Clinical Decision Support through Multi-Agent Systems and Ethical AI Governance

### v2

@@ III-G Quantitative Evaluation of Agent Transparency
Here we built the transparency assessment module [41] to evaluate the transparency of clinical prediction explanations by analyzing text responses for key transparency features. It calculates a transparency score by checking for the presence of five critical elements: explicit weights, monotonic relationships, feature importance, decision path explanations, and uncertainty quantification. The scoring process uses regular expression pattern matching to detect relevant terms in the response content. The final transparency score is normalized by dividing the raw score (number of detected features) by the maximum possible score (total number of transparency features), resulting in a value between 0 and 1 that quantifies how transparent and interpretable a clinical prediction explanation is.
@@ IV Results
The experimental results in Table I and Figure 2 demonstrate that the multi-agent system consistently outperforms the single-agent approach across all evaluated metrics over 150 unique patients. Each experiment was conducted across eight runs with approximately 150 patients per run, and the reported values represent the average performance to ensure consistency and robustness of evaluation. In terms of mortality prediction accuracy, the multi-agent model achieved a mean of 59%, while the single-agent model reached only 56%. This 3 percentage point improvement is consistent across multiple runs and represents enhancement in predictive capability in the high-stakes ICU environment. Additionally, the standard deviation for the multi-agent model is marginally higher, suggesting a bit more variability across runs.
@@
Our analysis indicates an enhancement in predicting Length of Stay (LOS) when using the multi-agent approach. In our study, the average prediction error drops to 4.37 days under the multi-agent strategy, compared to 5.82 days observed with the single-agent method—an improvement of roughly 25% in accuracy. This gain is important, given its direct influence on how ICU resources are allocated and care is planned. In addition, the multi-agent system registers a mean squared error of 35.49 as opposed to 48.13, along with a root mean squared error of 5.95 compared to 6.94, signifying that it delivers more stable and consistent predictions with fewer extreme fluctuations.
A particularly aspect of our findings is how the multi-agent approach manages to reduce the LOS prediction error. A mean error of 4.37 days, as opposed to 5.82 days from the single-agent method, illustrates a noteworthy improvement of 25%, which is critical in fine-tuning patient care. Moreover, the improved metrics—lower mean squared error (35.49 instead of 48.13) and root mean squared error (5.95 compared to 6.94)—further confirm that this system not only enhances accuracy but also offers more stable predictions across various patient groups.
Figure 3: Comparison of Multi-agent and Single-agent frameworks across two key metrics: Mortality Prediction Accuracy (top) and Length of Stay Mean Error (bottom).
Figure 3 reveals a clear analysis across the eight test runs we conducted. Looking at the top graph, you can see the Multi-agent system (blue line) consistently outperformed the Single-agent approach (green line) in predicting mortality. The Multi-agent accuracy ranges from about 56% to nearly 60%, while the Single-agent stays between 55%–57%. What’s notable not just that it performed better, but that this advantage held steady across every single run.
The LOS prediction results in the bottom graph show more improvement for multi-agent model. The blue line stays below the green throughout all runs, with errors around 4–4.7 days compared to the Single-agent’s 5.7–6 days. That gap - somewhere around a day and a half - might not sound huge until you consider what it means for real patients and hospital planning. Most systems like this show more ups and downs, but here the multi-agent setup maintained its edge consistently.
TABLE II: Comparison of transparency score of Multi-agent and Single-agent Models (Average over 8 runs)
@@
Table II provides a side-by-side comparison of average transparency scores from eight independent runs. The data shows that the single-agent approach scores an average of 86.21% in contrast to 85.50% for the multi-agent model. This suggests that, under our current evaluation criteria, both models perform similarly in terms of transparency with the single-agent design being marginally more interpretable. Despite the distributed nature of the multi-agent system, it maintains nearly equivalent transparency levels, indicating that our shared memory architecture effectively preserves reasoning traceability across multiple specialized agents.
Overall, both the multi-agent and single-agent models perform similarly in terms of predictive accuracy and error, with only slight differences that do not strongly favor one over the other. Interestingly, the single-agent model shows a marginal advantage, which might be linked to its avoidance of the complexities that come with managing interactions among multiple agents. Without effective coordination, a multi-agent system may not fully benefit from its distributed learning structure, making centralized processing by a single agent a more effective option under these circumstances. This finding highlights that, in the absence of well-optimized inter-agent dynamics, the simpler single-agent architecture can sometimes outperform a more complex system. Therefore, unless additional factors—such as scalability or deployment constraints—demand a multi-agent setup, both models are viable choices based on the results observed.
@@ V Conclusion
In our recent analysis, we found that the multi-agent system delivers more accurate predictions than the single-agent model. For example, it not only improves mortality prediction rates and shortens the forecasted length of stay—but it does so with comparable interpretability. Specifically, the single-agent model scores 86.21% for transparency, whereas the multi-agent system comes in at a very close 85.50%. This minimal difference suggests that our multi-agent architecture effectively maintains transparency despite the inherent complexity of coordinating decisions among several specialized agents.
This balance between performance and understandability has serious implications in the clinical arena. In scenarios where prediction accuracy is essential—such as in critical care decisions—the multi-agent approach clearly offers benefits. Conversely, when it is crucial to explain the reasoning behind predictions to build clinician trust, the higher transparency of the single-agent model is more attractive.
Looking ahead, our research will concentrate on boosting the clarity of the multi-agent framework without compromising its predictive strengths. By enhancing inter-agent communication protocols and incorporating more advanced explanation mechanisms, we aim to develop a system that marries the superior predictive power of multi-agent architectures with the interpretability required for safe and effective use in critical care.

### v3

@@ III-G Quantitative Evaluation of Agent Transparency
We built the transparency assessment process, which is similar to [32] evaluating the transparency of clinical prediction explanations by analyzing text responses for key features. The transparency score is calculated based on explainability, interpretability, and traceability in the clinical prediction framework. The explainability score shows how the system justifies decisions based on feature importance identification, concentration of critical factors, clear reasoning, and stakeholder explanations. In addition, the interpretability measures how humans understand the model’s workings based on reasoning processes, prediction predictability, complexity, and alternative scenario analysis. The traceability assesses documentation quality by tracking input data sources, data transformations, model development history, and decision process. The overall transparency score is calculated by these components, showing strong performance but it indicates needs for improving interpretability.
@@ IV Results
Our results shown in Table I and Figure 2 demonstrate that the MAS consistently outperforms the SAS approach across all evaluated metrics over 150 unique patients. Each experiment was conducted across eight runs with approximately 150 patients per run, and the reported values represent the average performance to ensure consistency and robustness of evaluation. In terms of mortality prediction accuracy, the MAS achieved a mean of 59%, while the SAS reached only 56%. While this 3 percentage point difference was consistent across multiple runs, further statistical analysis with larger samples would be necessary to establish clinical significance. Additionally, the standard deviation for the MAS is marginally higher, suggesting a bit more variability across runs.
@@
Our analysis indicates an enhancement in predicting LOS when using the MAS. In our study, the average prediction error drops to 4.37 days under the MAS, compared to 5.82 days observed with the SAS—an improvement of roughly 25% in accuracy. This gain is important, given its direct influence on how ICU resources are allocated and care is planned. In addition, the MAS registers a mean squared error of 35.49 as opposed to 48.13, along with a root mean squared error of 5.95 compared to 6.94, signifying that it delivers more stable and consistent predictions with fewer extreme fluctuations.
A particular aspect of our findings is how the MAS manages to reduce the LOS prediction error. A mean error of 4.37 days, as opposed to 5.82 days from the SAS, illustrates a noteworthy improvement of 25%, which is critical in fine-tuning patient care. Moreover, the improved metrics—lower mean squared error (35.49 instead of 48.13) and root mean squared error (5.95 compared to 6.94)—further confirm that this system not only enhances accuracy but also offers more stable predictions across various patient groups.
Figure 3: A comparison of MAS and SAS performance across two key metrics: Mortality Prediction Accuracy (top) and Length of Stay Mean Error (bottom). The diagram displays results from eight independent runs. For mortality prediction, higher accuracy indicates better performance, while lower mean error values (in days) represent more accurate predictions for LOS.
Figure 3 reveals a clear analysis across the eight test runs we conducted. Looking at the top graph, you can see the MAS (blue line) consistently outperformed the SAS (green line) in predicting mortality. The MAS accuracy ranges from about 56% to nearly 60%, while the SAS stays between 55%–57%. What’s notable not just that it performed better, but that this advantage held steady across every single run.
The LOS prediction results in the bottom graph show more improvement for MAS. The blue line stays below the green throughout all runs, with errors around 4–4.7 days compared to the Single-agent’s 5.7–6 days. That gap - somewhere around a day and a half - might not sound huge until you consider what it means for real patients and hospital planning. Most systems like this show more ups and downs, but here the MAS maintained its edge consistently.
TABLE II: Comparison of transparency score of MAS and SAS (Average over 8 runs)
@@
Table II provides a side-by-side comparison of average transparency scores from eight independent runs. The data shows that the SAS approach scores an average of 86.21% in contrast to 85.50% for the MAS. This suggests that, under our current evaluation criteria, both models perform similarly in terms of transparency with the SAS design being marginally more interpretable. Despite the distributed nature of the MAS, it maintains nearly equivalent transparency levels, indicating that our shared memory architecture effectively preserves reasoning traceability across multiple specialized agents.
Overall, the MAS outperforms SAS in terms of predictive accuracy and error. Interestingly, our results provide a different perspective from previous research [33] that suggests MAS may fail through inter-agent misalignment. The shared memory architecture within MAS and transparency assessment process can establish a trustworthy CDSS. This finding highlights that a well-optimized inter-agent dynamic is more beneficial than a simpler SAS. Although our MAS offers a robust system, we still need to consider whether additional factors—such as scalability or deployment constraints—demand a MAS, based on the results observed.
@@ V Conclusion
In this study, we found that the MAS performs better than the SAS for the mortality prediction accuracy (59% vs. 56%, respectively). Additionally, our results indicate that MAS has a shorter predicted LOS while maintaining a similar degree of transparency. Specifically, the MAS scores 85.5% for transparency, whereas the SAS shows a very close 86.21%. This subtle difference suggests that our MAS effectively maintains transparency despite the nature of complicated coordinating decisions among several specialized agents.
The balance between performance and explainability in decision support systems has significant implications for clinical patient care. In critical care scenarios where prediction accuracy is important, our MAS approach presents significant advantages. While maintaining high transparency, the MAS system improves predictive performance and addresses the critical needs for both reliable outcomes and interpretable decision processes that healthcare professionals can incorporate into their clinical judgment. On the other hand, when explaining prediction rationales to build clinician trust is essential, the higher transparency within Single-Agent Systems (SAS) still needs further investigation. Our research will focus on enhancing the clarity of the MAS while preserving its predictive advantages. Through improved inter-agent communication protocols and more sophisticated explanation mechanisms, we aim to develop a system that combines the predictive capabilities of MAS with the interpretability essential for safe and effective implementation in critical care settings.